ROTATIONALLY-RESOLVED SPECTRA OF THE Ã-X̃ SYSTEM OF C_{3} − NE AND C_{3} − AR VAN DER WAALS COMPLEXES

GUIQIU ZHANG, DEPARTMENT OF CHEMISTRY, SHANDONG TEACHER’S UNIVERSITY, PEOPLE’S REPUBLIC OF CHINA; JUN-MEI CHAO and \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECULAR SCIENCES, ACADEMIA SINICA, P.O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C. AND DEPARTMENT OF CHEMISTRY,NATIONAL TAIWAN UNIVERSITY, TAIPEI 107, TAIWAN, R. O. C..

A preliminary report of our rotational analysis of the Ã-X̃ system of C_{3} −Ne and C_{3} −Ar van der Waals (vdW) complexes will be given. The complexes were generated from 193-nm photolysis of a gas mixture containing allene and rare-gas atom under supersonic molecular beam condition. The laser-induced fluorescence spectra near the C_{3}, Ã-X̃, 2_{0}^{2−}K_{0}^{1} band were recorded with a resolution of 0.03cm^{−1}. Both a- and c-type transitions were observed. The rotational constants and geometry of these vdW complexes will be presented.